5. Recombination An individual with two homologous chromosomes with genotypes DeF and dEf generates the following 100 egg cells: DeF 35 Def 2 DEF 11 DEf 2 de 3 def 9 dEF 3 dEf 35 a) What is the recombination rate between the E locus and the Flocus? Between D and E? And between D and F? b) Which gene is in the middle? c Draw the homologous chromosomes, label the alleles, and draw the recombination event(s) that lead to the DEF gamete.
Answer #1

From the given information it is observed that -

a) The recombination rate between loci will be calculated from the formula

Recombination Rate = no.of recombinant progeny/ total of progeny x 100%

Recombination rate between E and F locus = 11+2/ 100 x 100% = 14%

Recombination rate between D and E locus = 11+2/ 100 x 100% = 13%

Recombination rate between D and F locus = 11+35/ 100 x 100% = 46%

b) The gene that is in the midle is the E.

c) Following is the recombination pattern that lead to teh DEF gamete